5 times the difference of a number and 3 is 4 less than four times the number. What is the number?
Answer by mananth(16946)

let number be x
difference between number and 3 = x-3
5 times = 5(x-3)
4 times the number = 4x
4 less
4x-4
5(x-3) = 4x-4
5x-15=4x-4
5x-4x= -4+15
x=11